What mass of hydrogen peroxide must decompose to produce 65.88 g of H2O


1
Expert's answer
2022-05-13T03:12:03-0400

"2H_2O_2\rightarrow 2H_2O + O_2"

"M(H_2O)=2x1+16=18 g\/mole"

"n(H_2O)=65.88\/18=3.66 mole"

"n(H_2O_2)=n(H_2O)=3.66 mole"

"M(H_2O_2)=2x1+2x16=34 g\/mole"

"m(H_2O_2)=34x3.66=124.44 g"